OpenStudy (anonymous):

Can someone help me with the next step of this problem: I am doing implicit differentiation: 3(2)(x^2+y^2)(2x+2yy')=100(y+xy')

OpenStudy (anonymous):

You just want to isolate y'?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. Let me type this out: Start by foiling out the left side, distributing the 6, and distributing the 100 on the right. \[12x^3+12y^3y'+12xy^2+12yx^2y'=100y+100xy'\] Then get all the y' terms on one side. \[12y^3y'+12yx^2y'-100xy'=100y-12x^3-12xy^2\] Then factor out a y': \[y'(12y^3+12yx^2-100x)=100y-12x^3-12xy^2\] Divide over: \[y'=\frac{100y-12x^3-12xy^2}{12y^3+12yx^2-100x}\] Then you can re-plug in your original y to get it complete in terms of x.

OpenStudy (anonymous):

Oh, so on the left side, we must multiply out (x^2+y^2)(2x+2yy')?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and then multiply the 6 in

OpenStudy (anonymous):

Or you can divide the 6 over. But then you have to carry around a 50/3 :/

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

No problem :)